OpenStudy (anonymous):

Help once more :) please... a circle has the equation 4(x-5)^2+4y^2=4 find the center (h,k) and radius r

OpenStudy (anonymous):

That's not a circle; is it 4y^2?

OpenStudy (anonymous):

Assuming it is 4y^2: Formula for a circle is (x-h)^2+(y-k)^2=r^2, so divide both sides by 4 in yours to get (x-5)^2+(y-0)^2=1, so it's center is (5,0), and radius is sqrt(1) = 1

OpenStudy (anonymous):

sorry i fixed it :)

OpenStudy (anonymous):

@channie2003 Ok, then yeah, my answer above should be good :)

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

@channie2003 Yup, glad to help. All I ask for in return is wishes of good luck on my AP test tomorrow ;)

OpenStudy (anonymous):

well I wish you the best of luck on your AP test!!!!

OpenStudy (anonymous):

@channie2003 Haha, thank you very much! :)